# Trusted Knowledge Extraction for Operations and Maintenance Intelligence

### v1

@@ 3.6.4 Relation Extraction Evaluation
We evaluate the RE output qualitatively on the same set of 100 records used to create gold standards for the other KE tasks. The relations extracted by an RE system may take several structures, including relational tables, XML files, and relational triples. For our evaluation, we use relational triples, since they are compatible with KG construction technologies such as Neo4j. We evaluate the triples generated for each document as if they were components of a KG, and thus use metrics defined for KGs. We report syntactic accuracy, semantic accuracy, and consistency based on definitions in Chapter 7 of Knowledge Graphs (Hogan et al. (2021)), and further described in D. We also report the total number of triples generated for the entire FAA dataset and the percent of records with generated triples.
@@ 4.1 Overview
As described in Section 3.6, we have quantitative evaluations for NER, CR, and NEL as well as qualitative scores for RE. Table 8 shows the main results. [^8] All items in parentheses following the name of the tool are model names unless otherwise stated. For ReFinED, wiki_w_nums is an abbreviation for the model, wikipedia_model_with_numbers, and wikipedia refers to the entity set used. More detailed results follow in the respective subsections, including precision, recall, and task-specific metrics.
@@ 4.2 Named Entity Recognition
As described in Section 3.6.1, we have strong-match and weak-match F1 scores for the NER tools which can be seen in the tables below. Table 9 shows the label-agnostic results from evaluation on UTFAA, and Tables 10, 11, 12, and 13 show the SemEval F1 scores on the benchmark-annotated datasets. Each tool is evaluated on the benchmark-annotated dataset that corresponds to the set of entities it is trained to recognize, which is denoted in brackets. Note that we also provide scores for PL-Marker, since it performs NER as an intermediate step in RE, and outputs its named entities in a readily available file.
@@ 4.5 Relation Extraction
As described in Section 3.6.4, and D, we qualitatively evaluated RE for syntactic accuracy, semantic accuracy, and consistency. These scores can be found in Table 17, which orders the tools by number of triples evaluated, and highlights high-scorers in bold. Note that the rightmost column “Docs w/ Predicted Trip” denotes the percentage of the 2748 OMIn dataset records for which the corresponding tool extracted one or more triples.
@@ 6 Conclusion
6.1 Key Contributions
We present the Operations and Maintenance Intelligence (OMIn) Dataset, with Version 1 curated initially from raw FAA Accident/Incident data. OMIn is curated for KE in operations and maintenance, featuring textual descriptions of maintenance incidents characterized by mentions of aircraft systems and domain-specific shorthand. We release the gold standards prepared for NER, CR, and NEL as part of OMIn. This baseline expands the portfolio in the aviation operations and maintenance domain, since it offers records on a variety of subject matters, long enough to provide context and valuable information for extraction. OMIn is the first open-source dataset curated for KE in the operation and maintenance domains. It also contains structured data, such as details of the aircraft, failure codes, and dates. The structured data can be used in future work alongside natural language text to develop an integrated and mutually validating KE approach. While OMIn is currently based on aviation maintenance incident data, this data has qualities common to many sets of records or logs in the operation and maintenance domains, making it a valuable baseline. By publicizing this dataset on our GitHub repository, along with a Zenodo DOI (Mealey et al. (2024)), we offer it to the community and invite collaboration toward a robust, open-source KE dataset for the domain. We will respond to GitHub Issues and release subsequent versions in partnership with the community.
Hand in hand with the OMIn benchmark dataset, we present a comprehensive evaluation of sixteen tools across the four stages of the KE workflow. This evaluation provides insights into their performance on real-world data, limitations, and potential for improvement in the maintenance domain. We identified the challenges faced by these tools when dealing with maintenance-specific data, highlighting the need for domain-specific training and adaptation to enhance their effectiveness and reliability.
6.2 Future Work
Future work can build on this study in three main directions: enhancing data and annotations, adapting LLMs to the maintenance domain, and integrating structured knowledge for more interpretable, robust systems.
First, improving data quality and gold standards can directly benefit KE performance. Pre-processing techniques, such as spellcheck and acronym expansion, may reduce ambiguity in maintenance texts. More importantly, designing domain-specific entity and relation labels, and expanding the gold standard dataset, will allow evaluators to better represent models’ performance on the unique characteristics of the maintenance and operations domains. This expansion of the gold standard dataset will also enable future work in fine-tuning the evaluated tools on the OMIn dataset to increase operational readiness.
Second, the emergence of LLMs offers new possibilities for KE workflows for the maintenance domain. Smaller LLMs can be fine-tuned on curated datasets such as the FAA corpus to help bridge vocabulary gaps in maintenance texts. Moreover, agentic workflows, where LLMs plan and orchestrate subtasks, can combine flexible reasoning with specialized tools. This hybrid architecture allows smaller, interpretable models such as those tested in this paper to act as domain-specific “modules” whose outputs are refined and contextualized by a more general LLM. When integrated into an agentic workflow, small, specialized models provide stable, interpretable building blocks for domain-specific tasks. Their outputs can be verified, refined, and contextualized by the LLM, ensuring both robustness and accountability.
Finally, future work can involve deepening the integration of KE workflows with structured knowledge resources. Ontologies and ontology design patterns offer reusable specifications of domain-specific concepts and relationships that can guide extraction and support inference. Incorporating knowledge bases such as Wikidata and constructing new maintenance-specific ontologies can improve knowledge extraction and decision support applications. The open release of our gold standard dataset lays the groundwork for continual curation of domain-annotated data as domain-specific knowledge structures are developed. We hope that the release of our dataset enables the community to explore both the incremental improvement of KE tools and the strategic incorporation of structured knowledge resources to advance the field toward more trustworthy and dynamic knowledge extraction systems for operations and maintenance.
@@ Appendix A CoNLL-2012 Format Pre-Processing
For CoNLL-2012, documents were organized in a tabular structure with each word on a separate line and more than 11 columns detailing the word’s semantic role. These columns include parts of speech, parse bits, predicate lemmas, speakers, and named entities. The parse bit links each word to a segment of the sentence’s Propbank-style parse tree, originally generated using a Charniak parser by the CoNLL-2012 developers. To recreate their process, we used the Charniak parser from Brown Laboratory for Linguistic Information Processing (BLLIP), along with the NLTK POS-tagger and spaCy EntityRecognizer. The CoNLL-2012 developers used the IdentifinderTM tool from BBN (Miller et al. (1998)) for NER; however, we used spaCy since it was more up-to-date. All records were uniformly labeled with ‘speaker1’ in the speaker field. We found through experimentation that the selected CoNLL-2012-based tools did not use the predicate-lemma and its associated columns, so we inserted placeholders in these fields. The OMIn data formatted for CoNLL-2012 is available in our repository [hidden for blind revision].
@@ C.2 OT Metrics Details
where $\text{jc}(c_{1},c_{2})$ denotes the distance between concepts $c_{1}$ and $c_{2}$, $\text{mss}(c_{1},c_{2})$ is the most specific subsumer of $c_{1}$ and $c_{2}$, and $p(c)$ is the probability of encountering an instance of concept $c$. In USC-ISI-I2 (2021), they use instance counts of a class to compute the probability $p(c)$ and normalize Jiang Conrath distance onto a $[0\ldots 1]$ similarity measure by dividing by the largest possible distance between $c_{1}$ and $c_{2}$ through the entity node (Q351201) in the ontology.
@@ D.4 Consistency
Consistency is the degree to which the set of output triples for each record/document is free of contradictions. Percent consistency is calculated via the expression: ($N_{triples}$ $N_{inconsistencies}$)/($N_{triples}$), where $N_{inconsistencies}$ is the number of triples such that if they were removed from the set of output triples, the remaining set would be consistent. For example, if there are 3 triples generated for a document and 2 of them contradict each other, there is 1 inconsistency since if one of the contradicting triples were removed, the remaining 2 would be consistent. In this case, it would receive a consistency score of 0.6667.

### v3

@@ 5.2.4 Relation Extraction Evaluation
We evaluate the RE output qualitatively on the same set of 100 records used to create gold standards for the other KE tasks. The relations extracted by an RE system may take several structures, including relational tables, XML files, and relational triples. For our evaluation, we use relational triples, since they are compatible with KG construction technologies such as Neo4j. To assess how well the generated triples support accurate KG construction, we use three KG evaluation metrics from Chapter 7 of Hogan et al. (2021), along with the total number of triples generated and the percentage of records containing triples. Our implementation methodology is detailed in  D.
  • Syntactic accuracy: measures grammatical agreement of triples.
  • Semantic accuracy: measures correctness of meaning.
  • Consistency: measures absence of contradictions across triples.
  • Number of Triples: total number of triples generated.
  • Percent Docs with Triple: percentage of records with one or more generated triples.
@@ 6.1 Overview
As described in Section 5, we have quantitative evaluations for NER, CR, and NEL as well as qualitative scores for RE. Table 7 shows the main results. [^7] All items in parentheses following the name of the tool are model names unless otherwise stated. For ReFinED, wiki_w_nums is an abbreviation for the model, wikipedia_model_with_numbers, and wikipedia refers to the entity set used. More detailed results follow in the respective subsections, including precision, recall, and task-specific metrics.
@@ 6.2 Named Entity Recognition
As described in Section 5.2.1, we have strong-match and weak-match F1 scores for the NER tools which can be seen in the tables below. Table 8 shows the label-agnostic results from evaluation on UTFAA, and Tables 9, 10, 11, and 12 show the SemEval F1 scores on the benchmark-annotated datasets. Each tool is evaluated on the benchmark-annotated dataset that corresponds to the set of entities it is trained to recognize, which is denoted in brackets. Note that we also provide scores for PL-Marker, since it performs NER as an intermediate step in RE, and outputs its named entities in a readily available file.
@@ 6.5 Relation Extraction
As described in Section 5.2.4, and D, we qualitatively evaluated RE for syntactic accuracy, semantic accuracy, and consistency. These scores can be found in Table 16, which orders the tools by number of triples evaluated, and highlights high-scorers in bold. Note that the rightmost column “% Docs w/ Predicted Trip” denotes the percentage of the 2748 OMIn dataset records for which the corresponding tool extracted one or more triples.
@@ 8 Conclusion
We introduced the Operations and Maintenance Intelligence (OMIn) dataset that provides a curated subset of FAA Accident/Incident narratives with gold standards for NER, CR, and NEL to support knowledge extraction in maintenance settings. OMIn pairs terse, shorthand-heavy text with structured fields (e.g., aircraft details, failure codes, dates) and, to our knowledge, constitutes the first open-source KE resource focused on operations and maintenance. Using OMIn, we conducted a zero-shot, on-premises evaluation of sixteen openly available tools across NER, CR, NEL, and RE. The results show that off-the-shelf systems struggle with the abbreviations, non-standard syntax, and short narratives typical of maintenance text: NER and NEL exhibit notably low recall, CR transfers more robustly, and RE methods display a precision–coverage trade-off. Together with our reproducibility observations, these findings suggest low technology readiness for immediate operational deployment without domain adaptation. OMIn v1 is derived from pilot/crew narratives rather than technician logs, and we do not provide a unified RE gold standard due to heterogeneous relation schemas; these are deliberate scope choices for a first baseline.
These outcomes directly motivate future work: normalization and acronym expansion, enriched domain-specific labels, and larger gold standards are the most immediate levers to raise recall and comparability. A shared maintenance ontology would enable a fair RE benchmark and support joint models that fuse text with OMIn’s structured fields. Resources will be released via our GitHub repository and an anonymized Zenodo DOI (withheld for review), and we invite community contributions to iterate OMIn and its benchmarks.
@@ Appendix A CoNLL-2012 Format Pre-Processing
For CoNLL-2012, documents were organized in a tabular structure with each word on a separate line and more than 11 columns detailing the word’s semantic role. These columns include parts of speech, parse bits, predicate lemmas, speakers, and named entities. The parse bit links each word to a segment of the sentence’s Propbank-style parse tree, originally generated using a Charniak parser by the CoNLL-2012 developers. To recreate their process, we used the Charniak parser from Brown Laboratory for Linguistic Information Processing (BLLIP), along with the NLTK POS-tagger and spaCy EntityRecognizer. The CoNLL-2012 developers used the IdentifinderTM tool from BBN (Miller et al. (1998)) for NER; however, we used spaCy since it was more up-to-date. All records were uniformly labeled with ‘speaker1’ in the speaker field. We found through experimentation that the selected CoNLL-2012-based tools did not use the predicate-lemma and its associated columns, so we inserted placeholders in these fields. The OMIn data formatted for CoNLL-2012 is available in our repository (Mealey et al. (2024b)).
@@ C.2 OT Metrics Details
where $\text{jc}(c_{1},c_{2})$ denotes the distance between concepts $c_{1}$ and $c_{2}$, $\text{mss}(c_{1},c_{2})$ is the most specific subsumer of $c_{1}$ and $c_{2}$, and $p(c)$ is the probability of encountering an instance of concept $c$. In USC-ISI-I2 (2021), they use instance counts of a class to compute the probability $p(c)$ and normalize Jiang Conrath distance onto a $0\ldots 1]$ similarity measure by dividing by the largest possible distance between $c_{1}$ and $c_{2}$ through the entity node (Q351201) in the ontology.
@@
QID Description Label Class JC
Q15766923 scientific journal Fuel 0.03 0.06
Q5507117 short-lived Bay Area post-hardcore musical… Fuel 0.00 0.00
Table 20: Example OT Measurements
@@ D.4 Consistency
Consistency is the degree to which the set of output triples for each record/document is free of contradictions. Percent consistency is calculated via the expression: ($N_{triples}$ - $N_{inconsistencies}$)/($N_{triples}$), where $N_{inconsistencies}$ is the number of triples such that if they were removed from the set of output triples, the remaining set would be consistent. For example, if there are 3 triples generated for a document and 2 of them contradict each other, there is 1 inconsistency since if one of the contradicting triples were removed, the remaining 2 would be consistent. In this case, it would receive a consistency score of 0.6667.